OpenStudy (anonymous):

Well... I need help again @jigglypuff314

OpenStudy (anonymous):

How do I solve a system by using a table?? what does it mean by a table

jigglypuff314 (jigglypuff314):

hmmm I think they might mean to plug in points as a table and see where they match? a table as in like | x | y1 | y2 | 0 2 0 1 4 3 2 6 6 <- then when the two y values match up at one x, that's the solution...

OpenStudy (anonymous):

It shows -0.2x - 0.3y = 0.2 and 0.4x - 0.3y = 1.4 Which lining it up , I can get rid of the y right ?

jigglypuff314 (jigglypuff314):

>.> they want you to do that by making a table...? erm... lemme see then...

OpenStudy (anonymous):

yesss .

jigglypuff314 (jigglypuff314):

| x | y1 | y2 | 0 0.666 3.5 0.1 0.2 to get the y1 and y2 values plug in each x value into the equations so if you plug in x=0 into the first equation, you would get 0.666 if you plug in x=0 into the second equation, you would get 3.5

OpenStudy (anonymous):

okay, now what do i do ?

jigglypuff314 (jigglypuff314):

keep on doing that for each x I gave so plug in x=0.1 into the first equation, then into the second equation and fill out the table

OpenStudy (anonymous):

okay... where do i stop??

jigglypuff314 (jigglypuff314):

when for the two y values you get, match for one x :P

OpenStudy (anonymous):

okay .. one second.

OpenStudy (anonymous):

so i would do it like -0.2(0.1)-0.3y=0.2? and solve for y ?

jigglypuff314 (jigglypuff314):

yep :)

jigglypuff314 (jigglypuff314):

whoops my mistake... try x=1 and x=2 instead

OpenStudy (anonymous):

okay one second!

jigglypuff314 (jigglypuff314):

sorry, I gtg to dinner now, will be back in less than 30min http://www.wolframalpha.com/input/?i=-0.2x+-+0.3y+%3D+0.2+and+0.4x+-+0.3y+%3D+1.4

OpenStudy (anonymous):

this is what i have so far after yours: x | y1 | y2 1 | 0 | -3.33 2 | -2 | -2

jigglypuff314 (jigglypuff314):

yep yep ^_^